OpenStudy (anonymous):

determine of the function is one-one and find the inverse. f(x)=x-9

OpenStudy (aum):

f(x) = x - 9 is the equation of a straight line. It is a one-to-one function because any horizontal line drawn will intersect the straight line at just one point. So it passes the horizontal line test make it a one-to-one function. Find the inverse: f(x) = x - 9 y = x - 9 solve for x. Add 9: y + 9 = x x = y + 9 Interchange x and y: y = x + 9 is the inverse.